A fireplace contains 46 bricks along its bottom row. If each row above decreases by 4 bricks, what is the function that shows the bricks on the 12th row?
1
Expert's answer
2019-08-11T07:27:23-0400

n-th term of the arithmetic progression is given by

"a_n=a_1+(n-1)d"

"a_1=46"

"d=-4"

"a_n=46+(n-1)(-4)=46-4n+4=50-4n"

n=12


"a_{12}=50-4\cdotp12=2"


Answer: 2